OpenStudy (anonymous):

PLEASE HELP find the solution to the following system y = x^2 - 2x -2 and y = 4x +5

OpenStudy (anonymous):

x^2-2x-2=4x+5 x^2-6x-7=0

OpenStudy (anonymous):

the answer has to be in coordinate form there should be 2 coordinates

OpenStudy (anonymous):

is that answer no solution?

OpenStudy (anonymous):

you can solve the equation given by Abdul Shabeer: \[x^2-6x-7 = 0\] using the abc-formula. Good luck!

OpenStudy (anonymous):

ok i guess i will do it that way

OpenStudy (anonymous):

but i need help finding the FULL answer